OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@paki

OpenStudy (paki):

solve these two equations for "x" and "y" and same with the below too... the similar values will show the equality... 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

ok gimme a sec

OpenStudy (anonymous):

for the first i got y=-3 and x=0

OpenStudy (paki):

hmmm... now perform the same for the below equations too... the similar set values are the required similar equations....

OpenStudy (anonymous):

for the 2nd one i got y=4.5 and x=-6

OpenStudy (paki):

are these values rite... mean your calculation is rite at all...?

OpenStudy (anonymous):

i wasnt talking bout the first two ansers i did it for these two 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

but idk if there rigth or not

OpenStudy (paki):

how you obtained (y=-3 and x=0)... can you show me your work here...

OpenStudy (anonymous):

\[6x+2y=-6~take~out~the~6x~then~solve~for~ 2y=-6\] thats what i do to solve for the certain variables

OpenStudy (paki):

NO....you did it wrong...

OpenStudy (anonymous):

thats what my teacher said to do but it might just be for something else i dont remember thats why i need help

OpenStudy (paki):

look at this... we will solve this like.... http://openstudy.com/study#/updates/52a0df3ee4b01a67794f53f3

OpenStudy (anonymous):

hey i brb

OpenStudy (anonymous):

ok that helps ty